OpenStudy (anonymous):

DEFINATE INTEGRAL: please assist - integral of 1/(x -sqrtx) from 4 to 9

OpenStudy (anonymous):

\[\int\limits_{4}^{9}\frac{ 1dx }{ x-\sqrt{x} }\]

OpenStudy (anonymous):

put\[\sqrt{x}=t,x=t^2,dx=2t~dt\] when x=4 \[t=\sqrt{4}=2\] when x=9, \[t=\sqrt{9}=3\] \[I=\int\limits_{t=2}^{3}\frac{ 2t~dt }{ t^2-t }=2\int\limits_{2}^{3}\frac{ dt }{t-1 }=?\]

OpenStudy (mathmale):

I've experimented with this problem and have seen that a simple substitution will greatly simplify this integration. Have you @mertich, considered a substitution yet?

OpenStudy (anonymous):

thanx thanx ! thanx

OpenStudy (anonymous):

I TRIED LET u =x, and i got du =1dx but as i further intergrated, i found that: ln(u -sqrt u) and inserting the limits i got ln6-ln2 = ln3, well its wrong because its ln4! I don't understand---

OpenStudy (anonymous):

@mathmale